A rectangular field is 80m by 330m. A rectangular barn 21m by 35m is built in the field. How much area is left over?
Answer by macston(5194)

A=remaining area
Area of field-area of barn=A
(80m*330m)-(21m*35m)=A
26400 sq m-735 sq m=25665 sq m ANSWER: The remaining area is 25665 square meters.